### v2

@@ 4.3 Scope and physical context
The present work is conceived as the $\mathcal{N}=1$ supersymmetric extension of our earlier analysis of $\mathfrak{sl}(3,\mathbb{R})$ JT–gravity Ozer:2025bpb . In that framework, asymptotic symmetry structures were studied from a bulk perspective, without recourse to Schwarzian–type boundary actions. The method focused on identifying residual gauge transformations and symmetry breaking patterns arising directly from the bulk gauge fields, providing a non–boundary formulation of the ASA.
In this supersymmetric extension, the same philosophy is maintained: we do not attempt to introduce a new effective boundary action, nor do we claim to formulate an alternative physical regime. Rather, we aim to generalize the bulk gauge–theoretic approach to the supersymmetric case, governed by the $\mathfrak{osp}(1|2)$ superalgebra.
The phrase “beyond the Schwarzian regime” is therefore not intended to suggest a new physical theory, but simply refers to a methodological shift away from boundary action–based analyses. This approach is rooted in the structure of gauge symmetries and their asymptotic realization, consistent with the strategy we established in the non–supersymmetric ${\tt S}{\tt L}(3,\mathbb{R})$ setting.
@@
In this sense, while our construction remains algebraic in scope, the extended bulk symmetry breaking mechanism within the $\mathfrak{osp}(1|2)$ framework may serve as a foundation for future developments involving boundary dynamics, supersymmetric observables, or holographic models that transcend the traditional Schwarzian description. We believe that this perspective may complement ongoing efforts to understand JT supergravity beyond its standard boundary effective action paradigm.
@@
From a physical standpoint, the dilaton-induced symmetry breaking mechanism analyzed in this work can be expected to influence observables in putative holographic dual theories. Specifically, the reduction of the full affine $\mathfrak{osp}(1|2)_{k}$ symmetry to its residual finite-dimensional subalgebra—together with the emergence of abelian commuting modes—modifies the spectrum of conserved charges at the boundary. This structure suggests a constrained class of boundary excitations, possibly leading to nontrivial selection rules or effective couplings in dual quantum mechanical models. While an explicit construction of the dual theory lies beyond the scope of this paper, we anticipate that the algebraic features derived here may provide structural constraints on supersymmetric SYK–like models or other one–dimensional superconformal systems. Further investigation of these physical implications remains an important direction for future work.
@@
We conclude this section by briefly commenting on the quantum aspects alluded to above. While the central extension of the superalgebra emerges naturally from the classical coadjoint orbit analysis, potential operator ordering ambiguities can affect the precise realization of the symmetry generators at the quantum level. Addressing such ambiguities may require a more systematic treatment—such as path integral quantization or BRST analysis—which lies beyond the scope of this work. Nonetheless, our current results offer a robust classical foundation for such future quantum investigations, particularly in supersymmetric generalizations of the Schwarzian theory.
@@
As a future direction, it would be interesting to explore quantum extensions of the symmetry breaking mechanisms presented here, especially in the context of super–Schwarzian dynamics or SYK–like boundary theories. The toy model introduced in Section 4.2 may also provide a useful testing ground for probing correlation functions and operator spectra in dual descriptions. We hope that our results provide a useful platform for further developments in two–dimensional supergravity and its holographic applications.
@@
Beyond the classical analysis, the incorporation of quantum aspects introduces subtleties that merit closer attention. One such feature is the appearance of central extensions, which in our setup arise naturally from the coadjoint orbit construction. In particular, the Kirillov–Kostant symplectic structure on the orbit allows for a well-defined symplectic form, from which the classical central term in the $\mathcal{N}=1$ superconformal algebra can be read off Grumiller:2017qao ; Ikeda:1993fh . However, when promoting this classical structure to the quantum level, further issues emerge.
Most notably, operator ordering ambiguities may affect the precise realization of the symmetry generators and their algebra. These ambiguities stem from the non-commutative nature of quantum operators, especially in systems with fermionic generators and nontrivial constraints, such as in the $\mathfrak{osp}(1|2)$ case. While the classical Poisson brackets yield a consistent algebra, the quantum commutators require careful normal ordering or regularization prescriptions to preserve associativity and supersymmetry.
To properly resolve these quantum-level challenges, several quantization approaches can be considered:
  • Coadjoint orbit quantization: constructs the Hilbert space by quantizing the orbit equipped with a symplectic form Grumiller:2017qao ; Ikeda:1993fh . This technique is geometrically transparent but faces challenges in the presence of superalgebraic structure and boundary conditions.
  • BRST quantization: systematically incorporates gauge symmetries and constraints into the quantum framework Ikeda:1993fh ; Grumiller:2013swa . It is particularly suitable for supergravity models and BF-type gauge theories but often trades off geometric intuition for algebraic rigor.
  • Path integral quantization: provides a functional integral formulation of the theory, enabling the study of partition functions and quantum correlation functions Engelsoy:2016xyb ; Cardenas:2018krd . For $\mathfrak{osp}(1|2)$ supergravity, this approach captures the boundary dynamics and fluctuations of super-Schwarzian modes.
While a detailed quantum treatment lies beyond the present scope, the formal framework established in this work offers a solid foundation for such investigations. In particular, the toy model and the algebraic symmetry breaking discussed in Section 4.2 provide concrete starting points for analyzing quantum deformations and the structure of dual one-dimensional superconformal systems.
@@ 5 Conclusions and Discussion
Two–dimensional gravity models, most notably JT gravity, where the dilaton plays an active, dynamical role, have become central in the study of low–dimensional holography and quantum gravity. In its $sl(2,\mathbb{R})$ formulation, JT gravity defines the asymptotic conditions of $\mathrm{AdS}_{2}$ spacetime and allows spacetime diffeomorphisms to be enhanced into an infinite–dimensional Virasoro symmetry. However, the presence of a time–dependent dilaton field induces a nontrivial deformation in this structure, dynamically breaking even the expected preservation of the ${\tt S}{\tt L}(2,\mathbb{R})$ subgroup within the full Virasoro algebra. This breakdown is directly linked to dilaton fluctuations at the boundary, which govern the emergent dynamics of physical quantities near the asymptotic region. Thus, the dilaton evolves from a mere geometric label into a core dynamical variable that reshapes the boundary metric structure itself.
This foundational mechanism is further developed in the context of an $\mathfrak{osp}(1|2)$ supergravity extension explored in this work, where symmetry breaking and symmetry extension are jointly realized within a unified gauge–theoretic formulation.
As a natural supersymmetric extension of classical $sl(2,\mathbb{R})$ JT gravity, the $\mathfrak{osp}(1|2)$ JT supergravity framework introduces a structurally richer theory that incorporates both higher–order gauge connections and a more intricate configuration of boundary symmetries. This is not merely a formal enlargement of the algebra via fermionic generators; rather, it entails a nontrivial extension of the dynamical phase space and the gauge sector itself.
Notably, the $\mathfrak{osp}(1|2)$ gauge fields exhibit a novel mechanism of symmetry breaking, one that does not ensure the full preservation of even the $\tt{O}\tt{S}p(1|2)$ subgroup, resulting in substantial modifications to the boundary dynamics. This extended algebraic framework plays a key role in exploring generalized holographic dualities, wherein extended symmetries affect thermodynamic quantities such as energy and entropy. The resulting theory provides fertile ground for investigating the interplay between supersymmetry, higher–spin fields, and low–dimensional holography.
In contrast to earlier formulations of $\mathfrak{osp}(1|2)$ dilaton supergravity, most of which rely on coadjoint orbit techniques or metric–based descriptions, our BF–theoretic approach provides a systematic derivation of the full $\mathcal{N}=1$ superconformal algebra. More importantly, it unveils novel dynamical effects induced by the dilaton fields that have no analogue within the $\mathfrak{sl}(2,\mathbb{R})$ framework. These effects arise naturally from the gauge–theoretic treatment and highlight the deeper physical implications of supersymmetric extensions, beyond their algebraic structure alone.
The primary distinction between $sl(2,\mathbb{R})$ and $\mathfrak{osp}(1|2)$–based JT gravity theories lies in the role of the dilaton field and its impact on the extension of asymptotic symmetry algebras. While the $sl(2,\mathbb{R})$ structure is sufficient to encode the isometric features of $\mathrm{AdS}_{2}$ geometry, the $\mathfrak{osp}(1|2)$ framework offers a richer mathematical landscape to investigate more intricate holographic dualities and the dynamical behavior of boundary physical quantities. In particular, $\mathfrak{osp}(1|2)$ provides a natural starting point for systematically extending the theory to higher–rank super Lie algebras and Chern–Simons–based formulations.
The findings of this study demonstrate that the asymptotic symmetry algebra is governed not only by the structural features of the gauge fields but also by the dynamical configurations of the dilaton field near the boundary. The dilaton is not an external matter sector, it is an intrinsic and structural part of the theory, taking values in the adjoint representation of the gauge algebra. Under suitable boundary conditions, its time–dependent components become dynamical and contribute to the asymptotic symmetry algebra in the form of an abelian ideal composed of mutually commuting modes. This extension is not the result of adding an external $U(1)$ symmetry but instead arises from internal degrees of freedom that become dynamically relevant at the boundary. Consequently, the commuting currents induced by the dilaton should be viewed as a natural substructure within the extended asymptotic symmetry algebra, consistent with a semi–direct sum structure. The boundary behavior of the dilaton therefore plays a pivotal role in enriching the algebra while preserving its internal coherence, revealing the significance of internal gauge dynamics in shaping the asymptotic structure of two–dimensional gravitational systems.
In conclusion, the presence of the dilaton field profoundly affects not only the boundary dynamics in both $sl(2,\mathbb{R})$ and $\mathfrak{osp}(1|2)$ JT gravity frameworks, but also the mechanisms of extension and breaking within their respective asymptotic symmetry algebras. These extended models serve as powerful tools for uncovering new symmetry structures tied to boundary conditions and significantly enrich the theoretical landscape of holographic duality. In particular, they provide a robust foundation for probing deeper into holographic constructions such as the AdS/CFT correspondence and the SYK model. Future investigations into higher–rank Lie group–based generalizations are expected to further advance our theoretical understanding in this domain.
As discussed in the Introduction, several recent studies have examined the connection between generalized JT gravity and higher–spin extensions, particularly in the context of deformed SYK models and enhanced symmetry structures. For instance, in Momeni:2024ygv , a deformed SYK model incorporating spin–$3$ operators is proposed as a potential holographic dual of generalized dilaton gravity. Our construction based on $\mathfrak{osp}(1|2)$, equipped with an explicit $\mathcal{N}=1$ superconformal symmetry and a controlled symmetry breaking framework, offers a concrete realization of this duality. In parallel, the works Momeni:2020tyt ; Momeni:2020zkx ; Momeni:2021jhk explore complementary approaches grounded in coadjoint orbit techniques, the quantization of extended superconformal algebras, and generalized boundary dynamics. While these efforts share similar motivations, our approach is distinguished by its BF–theoretic formulation, which systematically derives the asymptotic symmetry algebra and elucidates the role of dilaton-induced symmetry breaking. This perspective advances the theoretical understanding of how higher–spin structures can be coherently integrated into two–dimensional holographic frameworks.
In our model, the interplay between the dilaton and higher–spin fields at the boundary induces a breakdown of the asymptotic symmetry structure. The initially present infinite–dimensional $\mathcal{N}=1$ superconformal algebra becomes dynamically reduced to its $\tt{O}\tt{S}p(1|2)$ subgroup under the influence of the dilaton. This mechanism closely resembles Schwarzian–type symmetry breaking. However, beyond this complete reduction, the boundary fluctuations of the dilaton field can also support instances of symmetry extension. As a result, our framework exhibits a novel coexistence of both asymptotic symmetry breaking and symmetry enhancement. For a comprehensive summary of boundary theories and their corresponding symmetry structures, see Appendix B.
In conclusion, JT gravity, formulated as a two–dimensional BF theory, offers complementary theoretical frameworks for probing the geometric and physical origins of asymptotic symmetries and Schwarzian dynamics. The gauge symmetries inherent to BF theory, together with the algebraic structure of the target space, form the foundation of asymptotic symmetry algebras at the boundary. JT gravity, in this context, serves as a specific realization of such structures. In both formulations, the Schwarzian action emerges as an effective reduced model capturing the residual symmetries inherited from bulk gauge transformations. However, the dynamics governed by the Schwarzian action do not encompass the entirety of the gauge freedom, but rather reflect the behavior of a physically significant subset. Therefore, BF theory and JT gravity jointly provide a rich and unified approach to elucidating the full structure of boundary symmetries and their holographic implications.
@@
An open question remains as to whether the $\mathfrak{osp}(1|2)$ JT gravity framework developed in this work admits a well–defined quantum mechanical dual, possibly in the spirit of an extended SYK–like model involving spin-$\tfrac{3}{2}$ operators or deformed reparametrization modes. In our earlier $\mathfrak{sl}(3,\mathbb{R})$ analysis Ozer:2025bpb , we demonstrated how boundary dynamics can go beyond the classical Schwarzian regime. In a similar vein, the present study does not aim to reproduce the super–Schwarzian Cardenas:2018krd structure, but rather to explore boundary dynamics that transcend it. In this context, pursuing a consistent generalization based on extended $\mathfrak{osp}(\mathcal{N}|2)$ BF theories with $\mathcal{N}>1$ provides a promising direction for uncovering richer asymptotic structures and novel symmetry–breaking mechanisms in future work.
In addition, we have clarified how the affine–to–conformal transition may be interpreted as an infrared truncation, consistent with the dominance of pseudo–Goldstone boundary modes. We also outlined how coadjoint orbit, BRST, and path integral quantization methods may further illuminate the quantum structure of this framework. These observations suggest multiple pathways for extending the present work toward a more complete dual description.
@@
While our construction does not include an explicit boundary action, the structure of the residual $\mathfrak{osp}(1|2)$ symmetries obtained from the bulk gauge formulation suggests a connection to extended super–Schwarzian–type quantum mechanical models. We emphasize that our focus remains on the intrinsic gauge–theoretic realization of boundary symmetries, and not on deriving a holographic dual theory. Nevertheless, the algebraic structure revealed in our analysis may serve as a guiding framework for future efforts aiming to identify or construct boundary effective actions consistent with this extended supersymmetric ASA. Whether such a theory generalizes or deviates from the standard Schwarzian sector remains an open direction that we hope to explore in future work.
This residual symmetry structure thus serves as a partial bulk–boundary dictionary: while we do not construct an explicit boundary theory, the algebraic content of the $\mathfrak{osp}(1|2)$ BF formulation indicates a possible correspondence with a generalized super-Schwarzian boundary dynamics.

### v3

@@ 4.3 Scope and physical context
The present work develops the $\mathcal{N}=1$ supersymmetric extension of our earlier analysis of $\mathfrak{sl}(3,\mathbb{R})$ JT gravity Ozer:2025bpb . As in that framework, the asymptotic symmetry structure is derived directly from bulk gauge fields and residual gauge transformations, providing a boundary-independent formulation of the ASA. In the supersymmetric setting governed by the $\mathfrak{osp}(1|2)$ superalgebra, we maintain this bulk-centered perspective. The ASA is obtained intrinsically from the gauge-theoretic structure, without introducing an auxiliary effective boundary action. The emphasis is therefore placed on the structural origin of asymptotic symmetries within the bulk BF formulation. The phrase “beyond the Schwarzian regime” refers to this methodological standpoint. It signals a shift from boundary action–based reductions toward a direct analysis of gauge symmetries and their asymptotic realization. This perspective is consistent with the strategy established in the non-supersymmetric $\mathfrak{sl}(3,\mathbb{R})$ setting and extended here to the $\mathfrak{osp}(1|2)$ supergravity framework. In this sense, while our construction remains algebraic in scope, the extended bulk symmetry breaking mechanism within the $\mathfrak{osp}(1|2)$ framework may serve as a foundation for future developments involving boundary dynamics, supersymmetric observables, or holographic models that transcend the traditional Schwarzian description. We believe that this perspective may complement ongoing efforts to understand JT supergravity beyond its standard boundary effective action paradigm.
@@
From a physical standpoint, the dilaton-induced symmetry-selection mechanism analyzed in this work influences the spectrum of boundary observables in putative holographic dual theories. In the affine regime, the full $\mathfrak{osp}(1|2)_{k}$ current structure remains intact, while specific dilaton backgrounds dynamically select a residual subset of symmetry generators together with abelian commuting modes. This background-dependent selection modifies the spectrum of conserved charges and constrains the admissible class of boundary excitations. By contrast, the Drinfeld–Sokolov reduction discussed in Section 4.2 isolates the superconformal sector governing the infrared dynamics. In that regime, the pseudo-Goldstone mode associated with super-reparametrization symmetry becomes the dominant boundary degree of freedom, providing the standard low-energy description familiar from super-Schwarzian analyses. Although an explicit construction of the dual theory lies beyond the scope of this paper, the algebraic structures derived here provide structural constraints for supersymmetric SYK-like models and other one-dimensional superconformal systems. Further exploration of these implications remains an important direction for future work.
@@
We conclude this section by briefly addressing the quantum aspects associated with the classical analysis presented above. While the central extension of the superalgebra follows naturally from the classical coadjoint orbit structure, its precise operator realization at the quantum level requires a consistent treatment of ordering prescriptions. A systematic quantum formulation—such as path integral quantization or BRST analysis—lies beyond the present scope. Nevertheless, the classical results derived here establish a well-defined structural starting point for future investigations, particularly in the context of supersymmetric generalizations of the Schwarzian theory.
@@
Future work can extend the present analysis to quantum realizations of the symmetry-selection mechanisms identified here, particularly in the context of super-Schwarzian dynamics and SYK-like boundary theories. The toy model introduced in Section 4.2 offers a controlled setting for probing correlation functions and operator spectra in prospective dual descriptions. More broadly, the structural framework developed in this work provides a basis for further investigations of two-dimensional supergravity and its holographic applications.
@@
The classical analysis developed in the previous sections admits a natural quantum extension, but this step is accompanied by several structural subtleties. In our framework, central extensions arise geometrically from the coadjoint orbit construction. More precisely, the Kirillov–Kostant symplectic form defined on the orbit determines the classical central term of the $\mathcal{N}=1$ superconformal algebra in a coordinate-independent manner Grumiller:2017qao ; Ikeda:1993fh . In this sense, the central charge is not introduced ad hoc, but emerges from the underlying symplectic geometry of the reduced phase space.
Upon quantization, however, the transition from Poisson brackets to commutators introduces operator ordering ambiguities. These effects become particularly delicate in superalgebraic systems such as $\mathfrak{osp}(1|2)$, where fermionic generators and second-class constraints coexist. While the classical brackets close consistently, the quantum realization of the symmetry generators requires a careful choice of normal ordering prescription to preserve associativity, supersymmetry, and the graded Jacobi identities.
Several complementary quantization schemes may be employed to address these issues. Coadjoint orbit quantization provides a geometrically transparent construction of the Hilbert space by promoting the orbit symplectic form to a quantum commutator structure Grumiller:2017qao ; Ikeda:1993fh . In contrast, BRST quantization systematically incorporates gauge constraints and residual symmetries into a nilpotent cohomological framework Ikeda:1993fh ; Grumiller:2013swa , which is particularly well-suited to BF-type supergravity models. Finally, the path integral formulation captures quantum fluctuations around classical backgrounds and makes contact with the super-Schwarzian boundary theory and its partition function Engelsoy:2016xyb ; Cardenas:2018krd .
Although a full quantum analysis lies beyond the scope of the present work, the algebraic symmetry breaking mechanism and the Brown–Henneaux reduction discussed in Section 4.2 provide a controlled starting point for investigating quantum deformations of the boundary algebra. In particular, the affine-to-superconformal reduction suggests a natural framework for studying quantum corrections to the extended $\mathcal{N}=1$ symmetry and its dual one-dimensional superconformal dynamics.
@@ 5 Conclusions and Discussion
Two–dimensional gravity models, most notably JT gravity, provide a remarkably rich arena in which geometry, boundary dynamics, and holography intertwine. In the $\mathfrak{sl}(2,\mathbb{R})$ BF formulation, the theory fixes the asymptotic structure of AdS2 and promotes boundary diffeomorphisms to an infinite–dimensional Virasoro symmetry algebra. Our analysis shows that the realization of this symmetry structure is structurally sensitive to the dynamical role of the dilaton. Once the dilaton develops nontrivial time dependence at the boundary, the set of admissible residual gauge transformations is dynamically restricted: the expected ${\tt S}{\tt L}(2,\mathbb{R})$ subgroup embedded in the Virasoro algebra is no longer generically preserved as a stabilizer of the background configuration. This symmetry selection originates directly from boundary dilaton fluctuations and reflects the fact that the reduced phase space is controlled not only by the metric sector but also by dilaton degrees of freedom. Consequently, the dilaton is elevated from a geometric marker of curvature to a central dynamical variable that governs the boundary realization of asymptotic symmetries.
This foundational mechanism is further developed in the present work within an $\mathfrak{osp}(1|2)$ supergravity extension, where symmetry breaking and symmetry extension are realized simultaneously in a unified gauge–theoretic framework. In this supersymmetric setting, the dilaton supermultiplet not only constrains the admissible residual symmetries but also enriches the boundary algebraic structure, thereby linking affine and superconformal regimes within a single dynamical construction.
As a natural supersymmetric extension of classical $\mathfrak{sl}(2,\mathbb{R})$ JT gravity, the $\mathfrak{osp}(1|2)$ JT supergravity framework introduces a dynamically enlarged phase space through the inclusion of fermionic generators and the dilaton supermultiplet. This extension does not simply amount to a formal augmentation of the algebra; rather, it reorganizes the gauge sector and the boundary realization of symmetries while preserving the underlying affine structure. In particular, the supersymmetric formulation enriches the space of admissible residual symmetries and establishes a coherent link between affine and superconformal regimes within a unified gauge–theoretic description.
Notably, the $\mathfrak{osp}(1|2)$ gauge formulation reveals a background-dependent mechanism of symmetry selection, in which the realization of residual symmetries is dynamically constrained by the dilaton supermultiplet. Importantly, this does not deform the underlying affine structure, but rather restricts the subset of admissible generators compatible with a given boundary configuration. As a result, the boundary dynamics acquire additional structure relative to the purely bosonic case, while preserving the consistency of the gauge algebra. This extended framework provides a controlled setting for exploring generalized holographic correspondences in which supersymmetric symmetry realizations influence thermodynamic observables such as energy and entropy. In this sense, the theory offers a coherent arena for studying the interplay between supersymmetry and low-dimensional holography within a unified gauge-theoretic description.
In contrast to metric-based or purely coadjoint orbit formulations of $\mathfrak{osp}(1|2)$ dilaton supergravity, the BF-theoretic framework adopted here offers a unified gauge-theoretic derivation of the $\mathcal{N}=1$ superconformal algebra and its residual symmetry structure. Rather than postulating the boundary algebra a priori, it emerges directly from the analysis of admissible gauge transformations and boundary conditions. Within this formulation, the dynamical role of the dilaton supermultiplet leads to symmetry-selection effects that extend the structure encountered in the purely bosonic $\mathfrak{sl}(2,\mathbb{R})$ case. These effects arise naturally from the gauge-theoretic treatment and clarify how supersymmetric extensions modify the realization of boundary symmetries without deforming the underlying affine algebra.
The primary structural distinction between $\mathfrak{sl}(2,\mathbb{R})$ and $\mathfrak{osp}(1|2)$ JT gravity lies not in a mere algebraic enlargement, but in the dynamical role of the dilaton supermultiplet and its influence on the realization of asymptotic symmetries. While the $\mathfrak{sl}(2,\mathbb{R})$ formulation captures the isometric structure of $\mathrm{AdS}_{2}$ geometry and its Virasoro enhancement, the $\mathfrak{osp}(1|2)$ framework reorganizes the reduced phase space by incorporating fermionic generators and background-dependent symmetry selection mechanisms. This supersymmetric gauge-theoretic setting provides a controlled arena for analyzing richer boundary symmetry realizations and their holographic implications. In particular, it offers a natural starting point for systematic generalizations to higher superalgebras within the BF framework, in close analogy with Chern–Simons constructions of three-dimensional gravity.
The results of this work show that the asymptotic symmetry structure is determined not only by the gauge connection but also by the dynamical boundary configuration of the dilaton field. The dilaton is not an external matter sector; rather, it is an intrinsic component of the BF formulation, valued in the adjoint representation of the gauge algebra. Under suitable boundary conditions, time–dependent dilaton profiles dynamically select admissible residual symmetry realizations and generate additional commuting boundary modes. Importantly, this effect does not arise from introducing an external $U(1)$ sector, but from internal gauge degrees of freedom that become dynamically relevant near the boundary. The resulting commuting currents should therefore be understood as emerging from the intrinsic phase space structure, modifying the realization of the ASA without deforming its underlying affine structure. In this sense, the boundary behavior of the dilaton plays a decisive role in shaping the effective symmetry realization, highlighting how internal gauge dynamics govern the asymptotic structure of two–dimensional gravitational systems.
In conclusion, the dynamical role of the dilaton field plays a central part in shaping the boundary realization of asymptotic symmetries in both $\mathfrak{sl}(2,\mathbb{R})$ and $\mathfrak{osp}(1|2)$ JT gravity frameworks. Rather than deforming the underlying affine structures, the dilaton modifies the set of admissible residual symmetry realizations through background-dependent selection mechanisms. The supersymmetric extension studied here demonstrates how a unified gauge-theoretic formulation naturally accommodates both affine and superconformal regimes within a single coherent phase space description. This perspective clarifies the structural role of internal gauge dynamics in determining boundary observables and their symmetry properties. More broadly, such gauge-based constructions provide a systematic framework for exploring refined holographic correspondences, including connections to super-Schwarzian dynamics and SYK-like models. Extensions to higher-rank (super)algebras within the BF formalism offer a natural direction for future research in low-dimensional holography.
As discussed in the Introduction, recent studies have explored connections between generalized JT gravity, deformed SYK models, and extended symmetry structures in low-dimensional holography. In particular, Ref. Momeni:2024ygv analyzes deformed SYK-type models within the broader framework of generalized dilaton gravity. Within this context, the $\mathfrak{osp}(1|2)$ construction developed here provides a concrete gauge-theoretic realization of $\mathcal{N}=1$ superconformal symmetry together with a controlled mechanism of background-dependent symmetry selection. Complementary approaches based on coadjoint orbit techniques, quantization of extended superconformal algebras, and generalized boundary dynamics have been developed in Momeni:2020tyt ; Momeni:2020zkx ; Momeni:2021jhk . While sharing similar motivations, the present BF formulation differs in that the asymptotic symmetry structure is derived directly from admissible gauge transformations and boundary conditions, rather than postulated at the level of an effective boundary action. This unified gauge perspective clarifies how supersymmetric extensions and their associated boundary realizations can be consistently embedded within two-dimensional holographic frameworks.
In conclusion, JT gravity formulated within the two–dimensional BF framework provides a transparent gauge–theoretic description of the geometric origin of asymptotic symmetries and their boundary realizations. The bulk gauge symmetries, together with the algebraic structure of the target space, determine the admissible boundary transformations and the associated ASAs. Within this setting, the Schwarzian action arises as an effective description of a reduced sector of these symmetries, obtained after imposing suitable boundary conditions. While it captures the dynamics of the physically relevant reparametrization modes, it does not exhaust the full gauge structure present in the BF formulation. The BF perspective therefore clarifies how JT gravity encodes boundary symmetry realization in a systematic and internally consistent manner, providing a coherent framework for analyzing their holographic interpretation.
@@
An open question concerns whether the $\mathfrak{osp}(1|2)$ JT gravity framework developed here admits a well-defined quantum mechanical dual, potentially in the spirit of supersymmetric SYK-like models with extended boundary symmetry structures. In our earlier $\mathfrak{sl}(3,\mathbb{R})$ analysis Ozer:2025bpb , we showed that boundary dynamics can extend beyond the minimal Schwarzian sector. In a similar spirit, the present study does not attempt to reproduce the standard super-Schwarzian construction Cardenas:2018krd , but rather provides a gauge-theoretic setting in which alternative boundary realizations may consistently arise. Within this perspective, generalizations to extended $\mathfrak{osp}(\mathcal{N}|2)$ BF theories with $\mathcal{N}>1$ offer a natural direction for exploring richer asymptotic symmetry realizations in future work.
In addition, we have clarified how the transition from the affine to the superconformal regime may be understood as an infrared truncation, consistent with the dominance of pseudo–Goldstone boundary modes associated with super-reparametrization symmetry. This interpretation places the Schwarzian sector within a broader gauge-theoretic framework rather than treating it as a fundamental starting point. We have also outlined how coadjoint orbit, BRST, and path integral quantization approaches provide complementary perspectives for investigating the quantum realization of the boundary symmetry structure. Together, these observations indicate several coherent directions for developing a more systematic understanding of the quantum dual description.
@@
Although our construction does not introduce an explicit boundary action, the residual $\mathfrak{osp}(1|2)$ symmetry structure derived from the bulk gauge formulation naturally invites comparison with supersymmetric Schwarzian-type quantum mechanical models. The emphasis of the present work, however, is on the intrinsic gauge-theoretic realization of boundary symmetries rather than on constructing a specific holographic dual. The algebraic framework uncovered here may nevertheless provide structural constraints for prospective boundary effective descriptions consistent with the extended supersymmetric ASA. Determining whether such realizations reproduce, generalize, or depart from the standard super-Schwarzian sector remains an open and well-posed direction for future investigation.
The main contribution of this work is the systematic derivation of the asymptotic symmetry structure of two–dimensional $\mathcal{N}=1$ dilaton supergravity directly from the bulk BF formulation and its residual gauge transformations, without postulating an effective boundary action such as the super–Schwarzian. Within this framework, the dilaton supermultiplet does not act as a source of algebraic deformation, but rather as a dynamical selector that determines the admissible residual symmetry realizations in the asymptotic phase space. The resulting symmetry selection should therefore be understood as an intrinsic consequence of the bulk gauge structure, rather than as a restriction imposed a posteriori through holonomy or regularity conditions. In this way, our analysis complements existing treatments of boundary dynamics while providing a structurally refined gauge-theoretic perspective on the origin and realization of asymptotic symmetries in supersymmetric dilaton gravity.